what is the ratio of 17 1/2 inches to 2 1/3 feet? Not sure how to start the problem , do I convert the fraction into a mixed number or just work with the improper fraction?
Answer by richwmiller(17219)

17 1/2=35/2
2 1/3=7/3
35/2/7/3
35/2*3/7=(35/7)*3/2
5*3/2=15/2=7 1/2 or 7.5
you are confusing terms
17 1/2 is a mixed number
35/2 is an improper fraction